### v3

# vCache: Verified Semantic Prompt Caching

Abstract
Semantic caches return cached LLM-generated responses for semantically similar prompts to reduce inference latency and cost. They embed cached prompts and store them alongside their response in a vector database. Embedding similarity metrics assign a numerical score to quantify the similarity between a request and its nearest neighbor prompt from the cache. Existing systems use the same static similarity threshold across all requests to determine whether two prompts can share similar responses. However, we observe that static thresholds do not give formal correctness guarantees, can result in unexpected error rates, and lead to suboptimal cache hit rates. This paper proposes vCache, the first verified semantic cache with user-defined error rate guarantees. It employs an online learning algorithm to estimate an optimal threshold for each cached prompt, enabling reliable cache responses without additional training. Our experiments show that vCache consistently meets the specified error bounds while outperforming state-of-the-art static-threshold and fine-tuned embedding baselines. We release the vCache implementation and benchmarks to support future research.
1 Introduction
Semantic caches can reduce large language model (LLM) inference latency by up to 100× by reusing previously generated responses for semantically similar prompts (Bang, 2023). Unlike exact-match caching, these systems can return cached responses not only for identical prompts but also for prompts that are similar in meaning, even if the input text differs (Razi et al., 2024). A semantic cache operates as follows. The cache $C$ embeds every prompt request $x$ into a vector $\mathcal{E}(x)\in\mathbb{R}^{d}$ and retrieves the semantically most similar prompt $nn(x)$, alongside its response $r(nn(x))$, from a vector database (Pan et al., 2024). The cache measures similarity (e.g., cosine similarity) between two embeddings using $s(x)=sim\bigl{(}\mathcal{E}(x),\mathcal{E}(y)\bigr{)}\in[0,1]$. If no sufficiently similar prompt is found, the LLM generates a response and adds its embedding along with the response to the vector database in the cache. In this paper, we study the reliability of semantic caches in returning correct responses for semantically similar prompts.
To determine whether a new prompt is sufficiently close to an existing prompt in the cache, state-of-the-art semantic caches rely on a user-selected threshold $t\in[0,1]$ (Bang, 2023; Li et al., 2024; Dasgupta et al., 2025; Razi et al., 2024; Sudarsan and MasayaNishimaki, 2024). If $s(x)\geq t$, the system performs exploitation (cache hit) by returning the cached response $r(\mathrm{nn}(x))$. Otherwise, it performs exploration (cache miss) by querying the model for a new response $r(x)$. The cache adds $\mathrm{\mathcal{E}}(x)$ to the vector database, stores $r(x)$ in its metadata, and returns $r(x)$.
However, selecting an appropriate threshold $t$ is nontrivial. If the threshold $t$ is set too low, the system may treat unrelated prompts as similar, resulting in cache hits where the retrieved response $r(\mathrm{nn}(x))$ differs from the correct output $r(x)$. These false positives reduce response quality and compromise cache reliability. If $t$ is too high, the system may forgo correct cache hit opportunities and invoke the model unnecessarily  (Rekabsaz et al., 2017).
Existing systems use the same static similarity threshold across all requests. Users either use a predefined threshold (e.g., 0.8) or determine one by testing multiple values upfront (Dasgupta et al., 2025; Li et al., 2024; Dan Lepow, 2025; Razi et al., 2024; lit, 2025; Bang, 2023). This approach assumes that similarity correlates uniformly with correctness across all prompts and their embeddings. However, two prompts may be close in embedding space yet require different responses. Figure 2 illustrates that correct and incorrect cache hits have highly overlapping similarity distributions, suggesting that fixed thresholds are either unreliable or must be set extremely high to avoid errors, making them suboptimal. Another significant limitation of existing semantic caches is the lack of error-rate guarantees. While the latency benefits of caching are appealing, the risk of returning incorrect responses can outweigh those advantages. For widespread adoption, semantic caches must adhere to user-defined error rate tolerances.
We propose vCache, the first verified semantic cache with theoretical correctness guarantees. vCache learns a separate threshold (Figure 1) for each embedding in the cache, capturing the threshold variability observed in Figure 2. It requires no offline training, is agnostic to the underlying embedding model, and dynamically adapts its thresholds to the data distribution it encounters. As a consequence, vCache is robust to out-of-distribution inputs. To our knowledge, no prior work in semantic caching learns thresholds in an online manner.
We adopt a simple probabilistic framework to bound the error rate conditioned on a learned per-embedding threshold. When deploying vCache, the user specifies a maximum error rate bound, and the system maximizes the cache hit rate subject to this correctness constraint. Let $\mathbf{vCache}(x)$ denote the response returned by vCache. Let $\tau$ denote the exploration probability—a value monotonically decreasing in the likelihood of being correct—and calibrated so that the overall error rate remains below the user-specified bound (see Section 4). The decision rule modeling the probability of being correct for whether to exploit the cached response or explore an LLM inference is given by:
$\mathbf{vCache(x)}=\begin{cases}r(\mathrm{nn}(x))&\mathrm{Uniform}(0,1)>\tau,\\ \mathrm{LLM}(x)&\text{otherwise}.\end{cases}$ (1)
We evaluate the effectiveness of vCache in terms of correctness guarantees and overall performance. To assess generalizability, we compare vCache across two embedding models, two LLMs, and five datasets. We find that vCache consistently meets the error-rate bounds and outperforms static threshold baselines, even when using fine-tuned embeddings. Specifically, it achieves up to 2×2$\times$2 × higher cache hit rates and reduces error rates by up to 6×6$\times$6 ×. Our main contributions are as follows:
  1. 1.
    We propose vCache, the first semantic cache that enforces a user-defined correctness guarantee by bounding the error rate.
  2. 2.
    We introduce an online threshold learning algorithm that requires no prior supervised training, adapts to the observed data distribution, and is agnostic to the choice of embedding model.
  3. 3.
    We demonstrate that embedding-specific, dynamic thresholds improve decision quality. By learning a separate threshold per cached embedding, vCache achieves equal or better performance compared to both static thresholding and fine-tuned embeddings.
  4. 4.
    We introduce three benchmarks that reflect classification-style, browser search queries, and open-ended prompt distributions.
  5. 5.
    We release the vCache implementation [^1] https://github.com/vcache-project/vCache and three benchmarks  [^2] https://huggingface.co/vCache to support future research.
Figure 1: The static threshold in existing semantic caches enforces naive decision boundaries, resulting in either low cache hit or high error rates. vCache’s embedding-specific and dynamic thresholds learn decision boundaries to guarantee a user-defined maximum error rate. Gradient shading reflects decreasing confidence in correctness as similarity to the cached embedding decreases.
Figure 2: Results from 45k samples in the SemCacheClassification benchmark. Motivates the need for dynamic, embedding-specific thresholds. Top: Similarity values of correct and incorrect explorations exhibit highly overlapping distributions. Bottom: Optimal per-embedding thresholds vary substantially. See Appendix C.3 for details.
2 Related Work
LLM inference is both computationally expensive and latency-sensitive. To reduce these costs, semantic caches enable systems to return previously generated responses for not only identical, but also semantically similar prompts. Among existing systems, GPTCache is the most widely adopted open-source implementation, with over 7.5k GitHub stars (Bang, 2023). Several organizations and frameworks, including AWS (Razi et al., 2024), Microsoft (Dan Lepow, 2025), Redis-based caching in LiteLLM (lit, 2025), waLLMartCache (Dasgupta et al., 2025), and SCALM (Li et al., 2024), have adopted similar architectural patterns. A central limitation of existing systems is their reliance on a single global similarity threshold to determine whether a cached response should be returned to the user. This approach assumes that similarity correlates uniformly with correctness across all prompts and embeddings. However, as illustrated in Figure 2, the similarity distributions of correct and incorrect thresholds vary widely across prompts and embeddings. We describe the experimental setup underlying Figure 2 in Appendix C.3.
Semantic Cache Optimization
The threshold dilemma illustrated in Figure 2 can be addressed via two approaches: optimizing the embedding space or learning more effective thresholding strategies.
Embedding Fine-tuning: One way to improve caching accuracy is to reshape the embedding space so that prompts yielding the same response are close together, while those with different responses are well-separated, increasing the reliability of threshold-based decisions. Zhu et al. (2024) propose a distillation-based method that fine-tunes embeddings to better align embeddings with equivalent responses. A related challenge arises in image retrieval, where systems must determine whether a nearest neighbor corresponds to the correct target class. Zhang et al. (2023a) address this by introducing the Threshold-Consistent Margin loss, which enforces tighter intra-class cohesion and clearer inter-class separation by selectively penalizing negative pairs. However, they require supervised training, are limited to open-source embedding models, and can fail to generalize to out-of-distribution data at inference time (Hajipour et al., 2024). vCache’s online learning algorithm does not require training, is model-agnostic, and generalized to out-of-distribution data (see SemCacheClassification benchmark in Figure 5).
Threshold Optimization: Threshold optimization attempts to learn a decision boundary over existing embeddings without modifying the embedding model itself (Zhang et al., 2023b). To our knowledge, no prior work in semantic caching learns thresholds online at inference time. Yet, as shown in Figure 2, the optimal similarity threshold varies significantly across embeddings, motivating embedding-specific and online threshold estimation. Related ideas have been explored in pen-set recognition and incremental learning. For example, Rudd et al. (2017) propose the Extreme Value Machine (EVM), which models class boundaries using extreme value theory to support generalization to unseen categories. However, these methods do not enforce user-defined error rate guarantees. In contrast, we introduce the first online algorithm that estimates per-embedding thresholds for semantic caches while satisfying a user-defined error bound.
Semantic Cache Guarantees
Even with high-quality embeddings and carefully tuned thresholds, semantic caches remain inherently approximate. Unless a threshold of 1.0 is used (effectively restricting cache hits to exact prompt matches), there is always a risk of returning incorrect responses (Razi et al., 2024; Jim Allen Wallace, 2024). Despite this, existing semantic caching systems rely on fixed thresholds to decide whether to return a cached response (Dasgupta et al., 2025; Li et al., 2024; Dan Lepow, 2025; Razi et al., 2024; lit, 2025; Bang, 2023). As a result, they offer no formal guarantees on accuracy or error rates, making it difficult to justify their reliability in production environments. To address this, we propose vCache, the first semantic caching system that combines competitive performance with a user-defined correctness guarantee, offering interpretable and controllable reliability.
3 Overview of Semantic Caching
Setup
Let $\{x_{1},x_{2},\ldots,x_{n}\}$ be the set of all prompts inserted into the cache, in that order. Note that this set excludes prompts for which a cache hit was found and served. Let $\mathcal{D}$ denote all the data stored in the cache. For each prompt $x$ inserted into the cache, we store its vector embedding $\mathcal{E}(x)\in\mathbb{R}^{d}$, the true response $r(x)=\mathrm{LLM}(x)$ produced by the LLM, and optional additional metadata $\mathcal{O}(x)$. Given $\mathcal{E}(x)$, the cache retrieves the most similar prompt from the vector database with an approximate nearest neighbor search (Arya et al., 1998), where
$\mathrm{nn}(x)=\arg\max_{y\in C}sim\bigl{(}\mathcal{E}(x),\mathcal{E}(y)\bigr{)}.$ (2)
In vCache, for a prompt $x_{i}$, the metadata $\mathcal{O}(x_{i})$ stores similarity and response match information for all future prompts $x_{j}$ (with $j>i$) such that $\mathrm{nn}(x_{j})=x_{i}$. The exact workings of vCache are presented in Algorithm 1. The sets $\mathcal{D}$ and $\mathcal{O}$ can be represented as follows:
$\mathcal{D}=\Bigl{\{}\;\bigl{(}\mathcal{E}(x_{i}),\,r(x_{i}),\,\mathcal{O}(x_{i})\bigr{)}\;\Bigr{\}}_{i=0}^{n}\qquad\mathcal{O}(x_{i})=\Bigl{\{}\;\bigl{(}s(x_{j}),\,c(x_{j})\bigr{)}\;\bigl{|}\;nn(x_{j})=x_{i}\;\Bigr{\}}_{j=i+1}^{n}$ (3)
where $s(x)\in[0,1]$ is the similarity between $x$ and its nearest neighbor $\mathrm{nn}(x)$ and $c(x)$ indicates if the cached response of $\mathrm{nn}(x)$, $r(\mathrm{nn}(x))$, matches the true response $r(x)$.
$s(x)=sim(\mathcal{E}(x),\mathcal{E}(\mathrm{nn}(x)))\qquad c(x)=\begin{cases}1&\text{if }r\bigl{(}nn(x)\bigr{)}=r(x),\\ 0&\text{otherwise}.\end{cases}$ (4)
Algorithm
Given a prompt, say $x$, we first compute the embedding $\mathcal{E}(x)$ and find its nearest neighbor $\mathrm{nn}(x)$. The caching policy $\mathcal{P}$ then determines whether we should use the cached response for this prompt ($\mathrm{exploit}$) or run the LLM inference ($\mathrm{explore}$). In case we decide to exploit, $r(\mathrm{nn}(x))$ is returned. Otherwise, we run $r(x)=\mathrm{LLM}(x)$, compute $s(x)$ and $c(x)$, update the observations $\mathcal{O}(\mathrm{nn}(x))$ and add $x$ to the database $\mathcal{D}$ using,
$\mathcal{O}(\mathrm{nn}(x))=\mathcal{O}(\mathrm{nn}(x))\cup\{s(x),c(x)\},\qquad\mathcal{D}=\mathcal{D}\cup\{(\mathcal{E}(x),r(x),\emptyset)\}.$ (5)
The key challenge lies in designing a decision policy $\mathcal{P}(\ldots)$, as it directly impacts both the cache hit rate and the overall error rate.
Policy of existing systems ($\mathbf{\mathcal{P}_{gptCache}(s(x))}$)
In existing semantic caching systems, the decision function is implemented as a fixed threshold rule. Given a user-defined threshold $t$, the cache exploits if $s(x)\geq t$ by returning the cached response $r(\mathrm{nn}(x))$. Otherwise, it explores by invoking the model for a response. As discussed in Section 2, this approach lacks formal guarantees and does not adapt to variation in similarity value distributions.
Policy of vCache ($\mathbf{\mathcal{P}_{vCache}(s(x),\mathcal{O}(\mathrm{nn}(x)),\delta})$)
vCache replaces static thresholding with an embedding-specific decision function that respects a user-defined maximum error rate bound $\delta$. If the function returns $\mathrm{exploit}$, the cache has enough certainty in the correctness of the response, exploits its knowledge, and returns the cached answer $r(\mathrm{nn}(x))$ (Algorithm 1, L5). If the function returns $\mathrm{explore}$, the cache has insufficient certainty and explores. Section 4 provides all the details of the vCache policy.
4 vCache
Given a user-defined maximum error rate $\delta$, vCache maximizes the cache hit rate while ensuring the probability of correctness remains above $1-\delta$. Instead of relying on an unreliable static threshold or fine-tuned embeddings, vCache models probability distributions to make cache hit or miss decisions. The distributions are specific to each embedding in the cache and model the probability of correct cache hits for a given similarity value. To remain dataset agnostic and avoid costly offline training, vCache estimates these distributions online by selectively generating labels for uncertain similarity values. Since generating a label requires an LLM inference, equivalent in cost to not using a cache, vCache minimizes such inferences. We refer to such inferences as $\mathrm{explore}$ and classify them as a cache miss. For a given prompt $x$, the cached response $r(\mathrm{nn}(x))$ is considered uncertain when the observations $\mathcal{O}({nn(x))}$ do not provide sufficient evidence to determine whether the cached response is correct ($c(x)=1$). In contrast, if $\mathcal{O}({nn(x)})$ provides sufficient evidence, vCache proceeds with $\mathrm{exploit}$ by returning the cached response $r(\mathrm{nn}(x))$ without an LLM inference. The rest of this section formalizes these ideas and explains the vCache policy, $\mathbf{\mathcal{P}_{vCache}}$, in detail.
Algorithm 1 vCache Workflow
1:$e_{x}\leftarrow\mathcal{E}(x)$
2:$y\leftarrow\mathrm{nn}(x)$
3:$s(x)\leftarrow sim(e_{x},\mathcal{E}(y))$
4:if $\mathcal{P}(s,\mathcal{O}(y),\delta)=\text{$\mathrm{exploit}$}$ then
5:     return $r(y)$
6:else
7:     $r(x)\leftarrow\mathrm{LLM}(x)$
8:     $c(x)=\mathbf{1}(r(x)=r(y))$
9:     $\mathcal{O}(y)\leftarrow\mathcal{O}(y)\cup\{(s(x),c(x))\}$
10:     if $\lnot c(x)$ then
11:         $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathcal{E}(x),r(x),\emptyset)\}$
12:     end if
13:     return $r(x)$
14:end if
vCache workflow for deciding whether to exploit a cached response (cache hit) or explore an LLM inference (cache miss). The decision relies on $\mathbf{\mathcal{P}_{vCache}}$ (Section 4) and respects the error rate guarantee $\delta$.
Figure 3: Exploration probability for $emb_{42}$ and $emb_{128}$. Top: Observations are perfectly separable. Bottom: Observations are overlapping. vCache selects the optimal $\hat{t}$ and adjusts the exploration probability based on the user-defined $\delta=0.01$.
4.1 User Guarantee
One of the key features of vCache is that it takes a user-defined maximum error rate, $\delta$, and ensures that the overall performance of the cache remains within this error bound. Let $\mathbf{vCache}(x)$ denote the response returned by vCache, regardless of whether the decision was to $\mathrm{explore}$ or $\mathrm{exploit}$. Then, an error rate guarantee of $\delta$ implies:
Definition 4.1 (user-guarantee).
An error-rate guarantee of $\delta$ for vCache implies that the marginal probability of vCache returning the correct answer, given any arbitrary prompt $x$, is lower bounded by $(1-\delta)$. In other words,
$\mathbf{Pr}(\mathbf{vCache}(x)=r(x))\geq(1-\delta).$ (6)
To achieve the error guarantee, vCache probabilistically decides when to $\mathrm{explore}$ and when to $\mathrm{exploit}$. Let $\mathbf{Pr}_{\mathrm{explore}}(x|\mathcal{D})$ be the probability that, given a prompt $x$ and having accumulated data $\mathcal{D}$, vCache decides to $\mathrm{explore}$. Then, we can decompose the probability that vCache is correct as,
$\mathbf{Pr}(\mathbf{vCache}(x){=}r(x))=\mathbf{Pr}(\mathrm{explore}|x,\mathcal{D})+(1-\mathbf{Pr}(\mathrm{explore}|x,\mathcal{D}))\mathbf{Pr}(c(x)=1|x,\mathcal{D}).$ (7)
This expression reflects two disjoint events. First, vCache decides to explore with probability $\mathbf{Pr}(\mathrm{explore}|x,\mathcal{D})$, and in this case, the output $\textbf{vCache}(x)$ is same as $\mathrm{LLM}(x)$ by design. In the second case, the vCache decides to $\mathrm{exploit}$ with probability $(1-\mathbf{Pr}(\mathrm{explore}|x,\mathcal{D}))$ and in this case, the probability of vCache being correct is represented as $\mathbf{Pr}(c(x)=1|x,\mathcal{D})$ using notation from the previous section. To ensure error guarantees are maintained, we should have,
$\mathbf{Pr}(\mathrm{explore}|x,\mathcal{D})\geq\frac{(1-\delta)-\mathbf{Pr}(c(x)=1|x,\mathcal{D})}{1-\mathbf{Pr}(c(x)=1|x,\mathcal{D})}=\tau_{\mathrm{nn}(x)}(s(x)).$ (8)
To meet the guarantees, vCache models the $\mathbf{Pr}(c(x)=1|x,\mathcal{D})$. This inequality provides an actionable constraint: if the estimated probability of correctness from the cache is high, the system may exploit; if the estimate is low, the system must explore. As long as $\mathbf{Pr}(\mathrm{explore}|x,D)$ is larger than the $\tau_{\mathrm{nn}(x)}(s(x))$, the guarantees are achieved. Notation of $\tau$ is chosen to emphasize that it is a function over similarities and is specific to each embedding in the cache.
Since vCache can only estimate $\mathbf{Pr}(c(x)=1\mid x,\mathcal{D})$ based on a limited number of samples, it accounts for the uncertainty in the estimation by considering the confidence band of $\mathbf{Pr}(c(x)=1\mid x,\mathcal{D})$. The modeling details and the vCache policy are presented in the next subsection.
4.2 vCache Modeling
vCache imposes a sigmoid parametric model on the relationship between similarity and correctness. Specifically, for an arbitrary prompt $x$, the probability of correct cache hit is defined as
$\mathbf{Pr}(c(x)=1|x,\mathcal{D})=\mathcal{L}(s(x),t,\gamma)=\frac{1}{1+e^{-\gamma(s(x)-t)}},$ (9)
where $s(x)\in[0,1]$ is the similarity of $x$ with its near neighbour. $t\in[0,1]$ is an embedding-specific decision boundary parameter, and $\gamma>0$ is a parameter controlling the steepness of the function. The sigmoid form is well suited for this task: it induces a smooth and monotonic relationship between similarity and correctness probability and enables efficient maximum-likelihood estimation (MLE) of the threshold $t$ from labeled data. By fitting a continuous likelihood function rather than enforcing a hard threshold, vCache generalizes better from limited observations.
The MLE estimates for the parameters, say $\hat{t}_{\mathrm{nn}(x)}$ and $\hat{\gamma}_{\mathrm{nn}(x)}$, using all the meta-data $\mathcal{O}_{\mathrm{nn}(x)}$ solves the binary cross entropy loss,
$\hat{t}_{\mathrm{nn}(x)},\hat{\gamma}_{\mathrm{nn}(x)}=\arg\min_{t,\gamma}\mathbb{E}_{(s,c)\in\mathcal{O}_{nn(x)}}\left[\Big{(}c\cdot\log(\mathcal{L}(s,t,\gamma))\Big{)}+\Big{(}(1{-}c)\cdot\log(1-\mathcal{L}(s,t,\gamma))\Big{)}\right]$ (10)
Note that these parameters belong to a specific embedding in the cache (specifically $\mathrm{nn}(x)$).
Since these estimates are based on limited number of samples, estimating the true $\mathbf{Pr}(c(x){=}1|x,\mathcal{D})$, and thus the correct $\tau_{\mathrm{nn}(x)}(s(x))$ is not possible. To ensure we still achieve guarantees, vCache, instead computes a upper bound, say $\hat{\tau}$ for $\tau_{\mathrm{nn}(x)}(s(x))$ using pessimistic values for $t,\gamma$ from the $(1-\epsilon)$ confidence band for these points for some $\epsilon\in(0,1)$. Let these estimates be $t^{\prime}(\epsilon),\gamma^{\prime}(\epsilon)$. We compute $\hat{\tau}$ using,
$\hat{\tau}=\min_{\epsilon\in(0,1)}\frac{(1-\delta)-(1-\epsilon)\mathcal{L}(s(x),t^{\prime}(\epsilon),\gamma^{\prime}(\epsilon)))}{1-(1-\epsilon)\mathcal{L}(s(x),t^{\prime}(\epsilon),\gamma^{\prime}(\epsilon))}\geq\tau_{\mathrm{nn}(x)}(s(x)).$ (11)
The details of why $\hat{\tau}\geq\tau_{\mathrm{nn}(x)}(s(x))$ and how to obtain confidence bands for $t$ and $\gamma$ are explained in Appendix B. Once we have $\hat{\tau}$, we have to ensure that the probability of exploration is above this value (see Eq 8 ). This is achieved by sampling a uniform random variable $u\sim\mathrm{Uniform}(0,1)$. If $u\leq\hat{\tau}$, the vCache explores, i.e., runs the LLM model to obtain the correct response. Otherwise, it exploits the cache by returning $r(\mathrm{nn}(x))$. This randomized policy ensures that, in expectation, the system explores sufficiently often to meet the correctness guarantee while maximizing cache usage when reliability is high. The exact algorithm of how $\mathrm{explore}$ and $\mathrm{exploit}$ decisions are made are presented in Algorithm 2.
Algorithm 2 vCache Policy $\mathcal{P}_{vCache}\bigl{(}s(x),\,\mathcal{O}(\mathrm{nn}(x)),\,\delta\bigr{)}$
1:function $\mathcal{P}_{vCache}$($s(x),\,\mathcal{O}(\mathrm{nn}(x)),\,\delta$)
2:     $\hat{t},\hat{\gamma}\leftarrow\arg\min_{t,\gamma}\textrm{LogisticLoss}(t,\gamma,\mathcal{O})$
3:                                   $\triangleright$ i.e solve Eq 16
4:     $\tau\leftarrow\min_{\epsilon\in[0,1]}\mathcal{G}_{\tau}(x,\hat{t},\delta,\epsilon)$
5:     $u\sim\mathrm{Uniform}(0,1)$
6:     if $u\leq\tau$ then
7:         return $\mathrm{explore}$
8:     else
9:         return $\mathrm{exploit}$
10:     end if
11:end function
1:function $\mathcal{G}_{\tau}$($s,\,\hat{t},\,\hat{\gamma},\,\delta,\,\epsilon$)
2:     $t^{\prime},\gamma^{\prime}\leftarrow\phi^{-1}(\hat{t},\hat{\gamma},1-\epsilon)$
3:     $\alpha\leftarrow(1-\epsilon)\ \mathcal{L}(x,t^{\prime},\gamma^{\prime})$
4:     $\tau\leftarrow\dfrac{(1-\delta)-\alpha}{1-\alpha}$
5:     return $\tau$
6:end function
4.3 vCache Algorithm
To summarize, the final vCache algorithm works as follows. First, for each incoming prompt $x$, it retrieves its nearest cached embedding $y=\mathrm{nn}(x)$ and fits a logistic decision boundary $\hat{t}(y)$ using all labeled examples observed for $y$. It then computes the $\hat{\tau}$ using Eq 11 by iterating over different values of confidence $\epsilon$. Then we use a uniform random variable $u\sim Uniform[0,1]$ and $\mathrm{explore}$ if $u\leq\tau$ and $\mathrm{exploit}$ otherwise.
vCache makes the following assumptions,
  • The data $\mathcal{D}$ received by the cache is independently and identically drawn from the underlying distribution
  • The true probability of correctness of response match given similarity, i.e. $\mathbf{Pr}(c(x)=1|\mathcal{D},x)$ is well represented by the sigmoid family of functions (Eq 9).
Under these two assumptions, vCache policy ensures user-provided error-rate guarantees. This result is summarized in the theorem below.
Theorem 4.1.
Let $\delta\in(0,1)$ be the user-provided maximum error tolerance. Let $\mathcal{D},|\mathcal{D}|=n$ be the set of prompts seen by vCache at an arbitrary point in time. Then under the assumptions that prompts $\mathcal{D}$ are drawn i.i.d. from underlying distribution and sigmoid family of functions (defined in Eq 9) correctly model the true likelihood of correctness for each embedding, the probability of correct response from vCache for any arbitrary prompt $x$, executed in an online manner in accordance with Algorithm 2, is guaranteed to be greator than $1-\delta$. i.e.
$\mathbf{Pr}(\mathbf{vCache}(x)=r(x)|\mathcal{D})\geq(1-\delta)\forall x,n$ (12)
5 Evaluation
For our experiments, we use two popular embedding models – GteLargeENv1-5 (Zhang et al., 2024) and E5-large-v2 (Wang et al., 2022), and two LLM models for generating responses – Llama-3-8B-Instruct (Dubey et al., 2024) and GPT-4o-mini (gpt, 2024), representing both high-quality proprietary models and efficient open alternatives. We use the HNSW vector database (Malkov and Yashunin, 2018) with cosine similarity, a standard metric for comparing vector embeddings in semantic caching systems (Bang, 2023; Li et al., 2024; Dasgupta et al., 2025). All experiments are conducted on a machine running Ubuntu 24.04.2 LTS, equipped with an Intel Xeon Platinum 8570 CPU (for running cache) and an NVIDIA B200 GPU with 192 GB of memory (for running LLMs and embedding models).
Baselines
We use the following Cache-settings in our experiments
  • GPTCache (zilliztech, ): SOTA cache using a static threshold for all embeddings (Parameters: threshold).
  • GPTCache + Fine-tuned embedding (Zhu et al., 2024): Changing the embedding model in GPTCache. Embedding models are fine-tuned on data and method provided by (Zhu et al., 2024) for better separation of correct and incorrect matches (Parameters: threshold).
  • vCache: This is our proposed method (Parameters: error-rate guarantee $\delta$).
  • vCache + Fine-tuned embedding (Zhu et al., 2024): Same as vCache, but uses a fine-tuned embedding model (Parameters: error-rate guarantee $\delta$).
Figure 4: vCache meets the user-defined maximum error rate bound $\delta$ with steadily increasing cache hit rates (vCache is learning). GPTCache shows increasing error and hit rates, illustrating the unreliability of static thresholds. Static baselines use fixed thresholds (0.99 top, 0.86 bottom). See Figure 5 for a threshold vs. $\delta$ Pareto comparison.
Datasets
To the best of our knowledge, no open-source benchmark currently exists for evaluating semantic caches. We introduce and open-source three diverse benchmarks designed to reflect common caching scenarios. Appendix D provides the complete dataset and benchmark cards.
  • SemCacheLMArena: A randomly sampled subset of 60,000 queries from the LM-Arena human preference dataset (Chiang et al., 2024), containing open-ended, user-generated prompts. It models chatbot scenarios where user inputs vary semantically but can require similar responses.
  • SemCacheClassification: A benchmark of 45,000 prompts derived from three classification datasets (Saurabh Shahane, 2023; Talmor et al., 2018; Ni et al., 2019), spanning 28 distinct labels. It simulates environments where LLMs are embedded in SQL queries to perform classification tasks (Liu et al., 2024b).
  • SemCacheSearchQueries: A random subset of 150,000 real-world queries from the ORCAS dataset (Craswell et al., 2020), consisting of short, search-style questions. It reflects scenarios where LLMs are used in web browsers to answer recurring user queries (Wang et al., 2024; Liu et al., 2024a), and caching might reduce redundant inference.
Metrics
We measure the following metrics. (1) Error Rate: (Lower is better) defined as the $FP/n$. where $n$ is the total number of prompts and FP is the number of false positives. (2) Cache hit rate: (Higher is better) defined as $(TP+FP)/n$ where TP and FP are true positives and false positives, respectively. Together, TP and FP measure the total number of cache hits. We also show ROC curves Hoo et al. (2017). For the non-deterministic evaluation of vCache (Algorithm 2, Line 5), we compute 95% confidence intervals using Wallis binomial confidence bounds and contingency tests Wallis (2013).
Figure 5: Pareto comparison across a range of thresholds and $\delta$ values. Each point represents a full run on 60k (SemCacheLmArena) or 45k (SemCacheClassification) samples. vCache generally outperforms static-threshold baselines. While it may slightly underperform for small $\delta$, its stability under increasing sample sizes (Figure 4) indicates greater long-term reliability than static thresholds.
vCache respects user-defined error-rate requirements
We evaluate whether vCache satisfies the user-defined error rate $\delta$ while maintaining competitive performance. As shown in Figure 4 (left), vCache consistently meets the maximum error rate across $\delta$ values, with actual error remaining below the specified bound. The small gap between maximum error rate and observed error stems from the conservative $t^{\prime}$ estimation, which ensures robustness and can be further refined. Notably, as the error rate stabilizes, vCache continues to increase its cache hit rate (Figure 4, right), demonstrating effective learning over time. In contrast, GPTCache baselines show increasing error rates with sample size, reflecting the inherent limitations of fixed thresholds even as hit rates improve.
Dynamic and embedding-specific thresholds are superior to static thresholds
We investigate whether semantic caches benefit from using dynamic, embedding-specific thresholds rather than a single static threshold. To evaluate this, we compare vCache against static-threshold baselines by systematically varying the threshold values for GPTCache-based baselines and the maximum error rate ($\delta$) for vCache-based caches across a feasible range. Each point in Figure 5 represents a complete evaluation over the full benchmark dataset, enabling a direct Pareto comparison between vCache and all static-threshold configurations. We find that vCache achieves better ROC curves, higher cache-hit rates at a given error rate, and consequently, significantly lower average latency. Specifically, in the SemCacheLMArena benchmark (Figure 5 top), vCache achieves up to 6× lower error, 2× higher cache hit rates, and reduced average latency. In SemCacheClassification (Figure 5 bottom), vCache outperforms all baselines for maximum error rates above approximately 1.5%, where static methods either violate the error constraint or underutilize the cache. For error rate bounds below 1.5%, vCache adopts a more conservative policy. This is a result of vCache’s modeling strategy, which prioritizes correctness by increasing exploration in regions of uncertainty.
Note that for a given threshold, GPTCache error rates have not yet converged and exhibit an increasing trend. Thus, the data presented here shows GPTCache in a more favorable light than it would appear had the error rates fully converged.
SemCacheSearchQueries Benchmark
The results for this benchmark are presented in Appendix C.1. We observe similar trends to those shown in Figure 4. vCache consistently meets the error guarantees set by the user, while the error rates of GPTCache continue to increase with more samples. In contrast to the previous two datasets, GPTCache’s error rates appear far from converging for SemCacheSearchQueries, potentially due to the high diversity of the dataset.
6 Limitations
There are two main limitations of vCache. First, for responses longer than a few words, vCache uses the LLM-as-a-judge approach (Zheng et al., 2023) to assess response equivalence (Algorithm 1, L8), which requires an additional inference call. However, this can be executed asynchronously outside the critical path, so it does not impact latency. Moreover, since the output is a single token (e.g., "yes" or "no"), the computational overhead is minimal (Leviathan et al., 2023). Second, vCache relies on the assumptions of i.i.d. data and a sigmoid function family to represent the probability of correctness. If these assumptions are violated, the analysis may not hold. Nonetheless, both assumptions are natural and appear to model most use cases well, as supported by our experimental results. Finally, while semantic caches introduce embedding overhead compared to LLM inference, this is generally offset by latency gains from cache hits, making the trade-off worthwhile on average (see Appendix C.2).
7 Conclusion
We introduced vCache, a semantic cache that provides correctness guarantees by learning an optimal similarity threshold for each cached embedding online. This approach addresses the limitations of static thresholds and embedding fine-tuning, and ensures that the error rate remains below a user-specified bound. Our experiments demonstrate that vCache consistently satisfies this guarantee while outperforming existing methods. These results suggest that reliable, interpretable caching for LLMs is both practical and deployable.
Acknowledgement
The authors thank Sebastien Gautier-Jean-Jacques Beurnier and Siavash Ameli for their invaluable contributions. They also thank the Sky Lab at UC Berkeley for their support.
Appendix A Broader Impacts
This work contributes to the advancement of machine learning by improving the efficiency of large language model (LLM) inference. By reducing both computational cost and latency, our approach makes LLM-based systems more accessible to organizations and individuals with limited computational resources, thereby lowering the barrier to adoption. In addition, by decreasing the frequency of full LLM invocations, our method reduces overall compute demand and, consequently, the environmental impact associated with training and operating large-scale AI infrastructure.
Appendix B vCache Modeling Proof
We provide the proof for Theorem 4.1 in this section.
Recall we use the following notation in the paper,
  • $x$ : prompt under consideration
  • $\mathcal{D}$ : data inserted into the cache
  • $\tau_{\mathrm{nn}(x)}(s(x))$ : minimum probability of exploration associated with embedding $\mathrm{nn}(x)$ at similarity value $s(x)$
  • $\mathbf{Pr}_{c}=1-\delta$
Computing the exact probability $\mathbf{Pr}(c(x)=1\mid\mathcal{D},x)$ is expensive, so we derive a simpler upper bound. If we have an upper bound, then computing an upper bound for $\tau_{\mathrm{nn}(x)}(s(x))$ is straightforward.
Lemma B.1 (Upper-Bounding $\tau$).
if $\mathbf{Pr}(c(x)=1|\mathcal{D},x)\geq\alpha$, then,
$\tau_{\mathrm{nn}(x)}(s(x))\leq 1-\frac{\delta}{1-\alpha}$
Proof.
Rewrite $\tau$ to isolate the unknown probability:
$\begin{split}\tau_{\mathrm{nn}(x)}(s(x))&=\frac{(1-\delta)-\mathbf{Pr}(c(x){=}1|\mathcal{D},x)}{1-\mathbf{Pr}(c(x){=}1|\mathcal{D},x)}\\ &=1-\frac{\delta}{1-\mathbf{Pr}(c(x){=}1|\mathcal{D},x)}\end{split}$ (13)
Next, suppose we have a known lower bound
$\mathbf{Pr}\bigl{(}c(x)=1\mid\mathcal{D},x\bigr{)}\;\geq\;\alpha.$
Then
[…]
and since $\delta>0$, it follows that
$\frac{\delta}{1-\mathbf{Pr}\bigl{(}c(x)=1\mid\mathcal{D},x\bigr{)}}\;\geq\;\frac{\delta}{1-\alpha}.$ (14)
Hence, to guarantee the exploration probability meets $\tau$, it suffices to ensure
[…]
Lemma B.2 (Lower-Bounding Cache-Correctness Probability).
Given $\mathcal{D}$, let $\hat{t}_{\mathrm{nn}(x)}$ and $\hat{\gamma}_{\mathrm{nn}(x)}$ be the MLE estimates computed as,
$\hat{t}_{\mathrm{nn}(x)},\hat{\gamma}_{\mathrm{nn}(x)}=\arg\min_{t,\gamma}\mathbb{E}_{(s,c)\in\mathcal{O}_{nn(x)}}\left[\Big{(}c\cdot\log(\mathcal{L}(s,t,\gamma))\Big{)}+\Big{(}(1{-}c)\cdot\log(1-\mathcal{L}(s,t,\gamma))\Big{)}\right]$ (16)
Let $t^{*}$ and $\gamma^{*}$ be the true parameters such that $\mathcal{L}(s(x),t^{*},\gamma^{*})$ is the true probability of correct cache hits. Consider an arbitrary $\epsilon\in(0,1)$. Let $t^{\prime},\gamma^{\prime}$ be such that,
$\mathbf{Pr}(t^{*}>t^{\prime}||\gamma^{*}<\gamma^{\prime})<\epsilon$ (17)
Then,
$\mathbf{Pr}(c(x)=1|\mathcal{D},x)\geq(1-\epsilon)\mathcal{L}(s(x),t^{\prime},\gamma^{\prime})$ (18)
Proof.
By the law of total probability:
$\begin{split}&\mathbf{Pr}(c(x)=1|\mathcal{D},x)\\ &=\mathbf{Pr}(t^{*}>t^{\prime}||\gamma^{*}<\gamma^{\prime})\cdot\mathbf{Pr}(c(x)=1|\mathcal{D},x,(t^{*}>t^{\prime}||\gamma^{*}<\gamma^{\prime}))\\ &\qquad+\mathbf{Pr}(\neg(t^{*}>t^{\prime}||\gamma^{*}<\gamma^{\prime}))\cdot\mathbf{Pr}(c(x)=1|\mathcal{D},x,t^{*}\leq t^{\prime},\gamma^{*}\geq\gamma^{\prime})\\ &\geq\mathbf{Pr}(\neg(t^{*}>t^{\prime}||\gamma^{*}<\gamma^{\prime}))\cdot\mathbf{Pr}(c(x)=1|\mathcal{D},x,t^{*}\leq t^{\prime},\gamma^{*}\geq\gamma^{\prime})\\ &\geq\;(1-\epsilon)\,\cdot\mathbf{Pr}\bigl{(}c(x)=1\mid\mathcal{D},x,t^{*}\leq t^{\prime},\gamma^{*}\geq\gamma^{\prime}\bigr{)}.\\ \end{split}$ (19)
$\begin{split}&(1-\epsilon)\,\cdot\mathbf{Pr}\bigl{(}c(x)=1\mid\mathcal{D},x,t^{*}\leq t^{\prime},\gamma^{*}\geq\gamma^{\prime}\bigr{)},\\ &\geq(1-\epsilon)\int_{0}^{t^{\prime}}\int_{\gamma^{\prime}}^{\infty}\mathbf{Pr}(t^{*}=t,\gamma^{*}=\gamma|t\leq t^{\prime},\gamma\geq\gamma^{\prime})\;\;\mathbf{Pr}(c(x)=1|\mathcal{D},x,t^{*}=t,\gamma^{*}=\gamma)dtd\gamma\\ &\geq(1-\epsilon)\left(\int_{0}^{t^{\prime}}\int_{\gamma^{\prime}}^{\infty}\mathbf{Pr}(t^{*}=t,\gamma^{*}=\gamma|t\leq t^{\prime},\gamma\geq\gamma^{\prime})dtd\gamma\right)\cdot\ \inf_{t\leq t^{\prime},\gamma\geq\gamma^{\prime}}\mathbf{Pr}(c(x)=1|\mathcal{D},x,t)\\ &=(1-\epsilon)\,\inf_{t\leq t^{\prime},\gamma\geq\gamma^{\prime}}\mathbf{Pr}(c(x)=1|\mathcal{D},x,t)\\ &=(1-\epsilon)\,\inf_{t\leq t^{\prime},\gamma\geq\gamma^{\prime}}\mathcal{L}(x,t,\gamma)\\ &=(1-\epsilon)\cdot\mathcal{L}(x,t^{\prime},\gamma^{\prime})\end{split}$ (20)
since, $\mathcal{L}(x,t_{1},\gamma)<\mathcal{L}(x,t_{2},\gamma)$, $\forall x$. $t_{1}>t_{2}$ and $\mathcal{L}(x,t,\gamma_{1})<\mathcal{L}(x,t,\gamma_{2})$, $\forall x$. $\gamma_{1}<\gamma_{2}$
Combining these results lets us set
$\alpha=(1-\epsilon)\,\cdot\ \mathcal{L}(x,t^{\prime},\gamma^{\prime})\quad\implies\quad\tau^{\prime}(\epsilon)=1-\frac{\delta}{1-\alpha},$ (21)
and then use $\tau^{\prime}$. To find the best $\tau^{\prime}$ closest to the actual lower bound of exploration probability, we search for the minimum $\tau^{\prime}$ over the entire range of $\epsilon\in(0,1)$ Thus,
$\tau^{\prime}=\min_{\epsilon\in(0,1)}\left[1-\frac{\delta}{1-(1-\epsilon)\mathcal{L}(s(x),t^{\prime}(\epsilon),\gamma^{\prime}(\epsilon)}\right]$ (22)
Confidence Bound on Optimal Threshold
To find $t^{\prime},\gamma^{\prime}$ from the estimated $\hat{t},\hat{\gamma}$, we can use the confidence intervals by assuming a uniform prior on $t^{*}$ and $\gamma^{*}$. Since, under uniform prior the distributions of $\mathbf{Pr}(t^{*}|\hat{t})$ and $\mathbf{Pr}(\hat{t},t^{*})$ are the same,
$\begin{split}\mathbf{Pr}(t^{*}|\hat{t})=\frac{\mathbf{Pr}(\hat{t}|t^{*})\mathbf{Pr}(t^{*})}{\mathbf{Pr}(\hat{t})}\\ \mathbf{Pr}(t^{*}|\hat{t})\propto\mathbf{Pr}(\hat{t}|t^{*})\\ \mathbf{Pr}(t^{*}|\hat{t})=\mathbf{Pr}(\hat{t}|t^{*})\end{split}$ (23)
Thus we can obtain $t^{\prime},\gamma^{\prime}$ using CDF of $\mathbf{Pr}(\hat{t},\gamma|t^{*},\gamma^{*})$
In our experiments, we only use confidence intervals for $t$, i.e., we use the $t^{\prime}$ parameter to adjust the likelihood. We estimate and use the point estimate $\hat{\gamma}$ for $\gamma$.
Appendix C Evaluation Results
C.1 SemCacheSearchQueries Benchmark
We discuss the SemCacheSearchQueries benchmark, focusing on understanding the limitations of static-threshold caching. We highlight why fixed thresholds fail to maintain reliable error guarantees at scale and how vCache addresses this issue through dynamic, embedding-specific thresholding.
Figure 6: vCache meets the user-defined maximum error rate bound $\delta$ with steadily increasing cache hit rates (vCache is learning). GPTCache shows increasing error and hit rates, illustrating the unreliability of static thresholds. The static baseline uses a fixed threshold of 0.985. See Figure 7 for a threshold vs. $\delta$ Pareto comparison.
vCache respects user-defined error-rate requirements
We evaluate whether vCache satisfies the user-defined error rate $\delta$ while maintaining competitive performance. As shown in Figure 6 (left), vCache consistently remains below the specified error bound across all tested $\delta$ values. Moreover, as the error rate stabilizes, vCache continues to improve its cache hit rate (Figure 4, right). In contrast, GPTCache exhibits increasing error rates as the sample size grows, despite improved hit rates. This trend reflects a fundamental limitation of static thresholds: maintaining a bounded error rate requires continuously increasing the threshold. Over time, no static threshold below 1.0 may suffice to satisfy a strict error constraint, making such systems difficult to tune and unreliable at scale.
Figure 7: Pareto comparison across a range of thresholds and $\delta$ values on 150k samples from the SemCacheSearchQueries benchmark. While vCache slightly underperforms static baselines at this scale, its error rate remains bounded. In contrast, static-threshold methods like GPTCache require increasing thresholds to maintain low error, shifting the Pareto curve upward over time. This highlights the limitations of static thresholds in sustaining error-rate guarantees as sample size grows.
Dynamic and embedding-specific thresholds are superior to static thresholds
We evaluate whether dynamic, embedding-specific thresholds yield better long-term performance than static thresholding. To this end, we compare vCache against GPTCache by varying static similarity thresholds and vCache’s maximum error rate bound $\delta$. Each point in Figure 7 reflects a full evaluation over the SemCacheSearchQueries benchmark, enabling a direct Pareto comparison. Static-threshold configurations achieve higher cache hit rates and lower latencies than vCache under equivalent evaluation settings. However, this performance comes at the cost of increasing error as the sample size grows, as shown in Figure 6. Because the threshold remains fixed, both hit rate and error rate rise together. As a result, the GPTCache curve in the middle plot (cache hit vs. error rate) is expected to shift up and to the right, while the curve in the right plot (error rate vs. latency) shifts up and to the left. This trend implies that no static threshold below 1.0 can maintain a bounded error rate as prompt diversity increases. In contrast, vCache learns its threshold online and per embedding, allowing it to enforce the error constraint while gradually improving cache hit rate. While this results in a more conservative reuse policy initially, vCache offers greater reliability under error-sensitive deployment conditions.
C.2 Embedding Generation Overhead
Semantic caches incur additional latency due to embedding computation, which must be evaluated in relation to the cost of LLM inference. To quantify this overhead, we compare the embedding latencies of four models—including fine-tuned variants—to the inference latency of Llama3.1-8B.
Figure 8: Embedding computation latency distributions across models, shown as 95th percentile whisker plots. GTE_FT = GteLargeENv1-5 [Zhang et al., 2024] fine-tuned [Zhu et al., 2024]. E5 Large v2 FT = E5-large-v2 [Wang et al., 2022] fine-tuned [Zhu et al., 2024]
All experiments are conducted on a system with an Intel Xeon Platinum 8570 CPU and an NVIDIA B200 GPU with 192 GB of memory. The results show that embedding computation is significantly faster than model inference, confirming its applicability in caching pipelines.
C.3 Threshold Dilemma
To analyze the relationship between similarity scores $s(x)$ and cache correctness $c(x)$, we conduct an experiment on the 45,000 entries of the SemCacheClassification benchmark using an error tolerance of $\delta=0.02$ (2%). For each cached embedding $x_{i}$, we record the observations $\mathcal{O}(x_{i})$ and the empirically estimated optimal threshold $\hat{t}_{x_{i}}$.
In the top plot of Figure 2, we separate all observations into two sets: one where $c(x)=1$ (correct cache hit) and another where $c(x)=0$ (incorrect hit). We plot the kernel density functions (KDF) to visualize the distribution of correct and incorrect observations. The result shows that correct and incorrect observations are nearly indistinguishable in similarity space, with substantial overlap and similar means (0.84 vs. 0.85). This illustrates that one single similarity threshold is not a reliable decision boundary.
The bottom plot shows a histogram of the optimal threshold values $\hat{t}$ computed per embedding. The thresholds span a range, from 0.71 to 1.0, indicating that no single similarity threshold can suffice across embeddings. A threshold set too low increases the risk of incorrect cache hits; a threshold set too high limits cache hits. Together, these results motivate the need for embedding-specific and dynamically learned thresholds to ensure interpretable and reliable performance.
Appendix D Benchmark Creation
To the best of our knowledge, no open-source benchmarks currently exist for evaluating the performance and applicability of semantic caching systems. To address this gap, we construct and release three diverse benchmarks [^3] https://huggingface.co/vCache, each designed to reflect a distinct real-world use case: classification tasks, conversational chatbots, and search engines. This section describes the motivation and construction process behind each benchmark.
D.1 SemCacheClassification Benchmark
The SemCacheClassification benchmark is designed to evaluate semantic caching in structured classification settings, such as those found in modern database environments [Liu et al., 2024b]. Several database systems—including Databricks, Snowflake, and AWS—have introduced LLM-based extensions to SQL via User-Defined Functions (UDFs), enabling capabilities beyond traditional SQL semantics [Liu et al., 2024b]. However, LLM-based UDFs are inherently non-deterministic, execute row-by-row, and pose integration challenges for parallelization and query optimization [Franz et al., 2024]. When table rows contain semantically similar values, semantic caching can reduce LLM invocation frequency, improving both latency and cost. This benchmark captures such use cases, where small variations in phrasing should not require repeated inference.
The benchmark consists of 45,000 short-form prompts with a fixed output label space. Each example follows a prompt–response format, where the prompt expresses a classification query and the expected response is a one-word label. The benchmark combines three diverse classification datasets: CommonsenseQA [Talmor et al., 2018], Ecommerce Categorization [Saurabh Shahane, 2023], and Amazon Instant Video Review [Ni et al., 2019]. Sample prompts and response formats from each dataset are shown below, and Table 1 summarizes label distributions across the benchmark.
A sample entry from the Ecommerce Categorization [Saurabh Shahane, 2023] dataset:
"prompt": "Which category does the text belong to? Text: <text>",
"output_format": "Answer with Books, Electronics, Household, or Clothing & Accessories only"
A sample entry from the Commonsense QA Talmor et al. [2018] dataset:
"prompt": "What is the main subject of the following question? Question: <question>",
"output_format": "Answer with only one of the words of this set: [people, potato, competing, , snake, lizard, food, car, water, student, crab, children, killing, animals, ficus, horse, fox, cat, weasel, shark, person, human]"
A sample entry from the Amazon Instant Video Review Ni et al. [2019] dataset:
"prompt": "Is this review friendly? Review: <review>",
"output_format": "Answer with yes or no only"
The benchmark enables controlled evaluation of semantic caching strategies, especially in cases where small changes in input phrasing must still map to the same output class. Its fixed label format makes it particularly useful for evaluating systems like vCache, which rely on correctness guarantees under threshold uncertainty. Table 1 summarizes the label distributions for the three subtasks in the SemCacheClassification benchmark.
Response Count
person 2,806
people 625
human 400
competing 272
animals 225
food 202
car 125
water 100
student 79
children 32
killing 27
horse 24
lizard 13
potato 13
fox 11
cat 10
ficus 10
weasel 8
shark 8
crab 7
snake 3
(a) Commonsense QA.
Response Count
Books 6,000
Clothing 6,000
Electronics 6,000
Household 2,000
(b) Ecommerce.
Response Count
yes 10,000
no 10,000
(c) Amazon Instant Video.
Table 1: Response distribution across three datasets that form the SemCacheClassification benchmark.
D.2 SemCacheLMArena Benchmark
The SemCacheLMArena benchmark is designed to evaluate semantic caching in chatbot environments, where users may issue semantically similar prompts with different phrasing. In such settings, caches must generalize across diverse surface forms while maintaining response correctness. This benchmark captures these challenges by grouping semantically similar user inputs and testing whether caching systems can reuse responses accurately.
To construct the benchmark, we use the LM-Arena human preference dataset [Zheng et al., 2023], which contains 100,000 real-world user queries. We randomly sample 3,500 distinct prompts, each of which defines a class. For each class, we generate between 1 and 23 semantically similar variants using GPT-4.1-nano, resulting in a total of 60,000 prompts. A class ID is assigned to each prompt to evaluate caching correctness: a cache hit is considered correct if the retrieved response belongs to the same class as the query. Figure 9 shows the distribution of class sizes (number of prompts belonging to a class), confirming broad variability in prompt paraphrasing. To support model-agnostic evaluation, we generate responses for all prompts using GPT-4.1-nano and GPT-4o-mini. The corresponding response length distributions are shown in Figure 10.
Figure 9: Distribution of class sizes in the SemCacheLMArena benchmark. Each class corresponds to a unique user prompt and contains 2–24 semantically similar variants.
Figure 10: Response length histogram for GPT-4.1-nano and GPT-4o-mini on the SemCacheLMArena benchmark. Median and mean lengths are shown for each model.
D.3 SemCacheSearchQueries Benchmark
The SemCacheSearchQueries benchmark is designed to evaluate semantic caching in open-domain search applications. Large-scale search engines, such as Google, increasingly rely on LLMs to generate direct answers to natural language queries [Liu et al., 2024a, Wang et al., 2024]. While this improves user experience, it introduces significant latency and cost, particularly at the scale of millions of daily queries. Many queries issued to search engines are paraphrased variations of earlier inputs, making semantic caching a natural fit for reducing redundant LLM inference in this setting.
The benchmark is constructed from a filtered subset of the ORCAS dataset [Craswell et al., 2020], containing real-world search engine queries. We begin by sampling 500,000 queries and embedding each using the gte-large-en-v1.5 embedding model [Zhang et al., 2024]. We then apply $k$-means clustering to group similar queries and retain the largest clusters, resulting in 150,000 entries. Within each cluster, we apply a union-find algorithm guided by an LLM-based judge (GPT-4.1-nano) to determine whether query pairs yield the same response. Sub-clusters identified in this step define the equivalence classes used for caching evaluation. Figure 11 summarizes the benchmark properties, including class size distribution, frequent query terms, and statistics on the number of queries per class.
Figure 11: Descriptive statistics for the SemCacheSearchQueries benchmark. Top left: Most classes contain exactly two queries. Top right: The word best appears in over 16,000 of the 150,000 queries. Bottom left: The largest class contains 90 semantically equivalent queries. Bottom right: 59.3% of all classes contain more than eight queries, indicating substantial intra-class variability.